# A Taxonomy of Questions for Critical Reflection in Machine-Assisted Decision-Making

### v2

@@ 2.5. The Socratic Questioning Method
A common technique to facilitate reflection is to ask questions (Osmond and Darlington, 2005). Moreover, questions encourage independent thinking, which can increase confidence in one’s own decision-making, i.e., decision self-efficacy, which in turn can affect the motivation and well-being of the decision-maker (Ryan and Deci, 2000). Findings from educational research suggest that motivation functions as an antecedent to cognitive engagement and also increases it (Rotgans and Schmidt, 2011; Blumenfeld et al., 2006; Singh et al., 2022). Questions thus seem to promote reflection, which can increase decision self-efficacy and motivation, which in turn can increase cognitive engagement.
@@
To formulate relevant and effective questions, we want to point out two works, which we also use for our purposes. First, Bloom’s taxonomy for educational objectives (Krathwohl, 2002; Anderson and Krathwohl, 2001) categorises cognitive processes into six levels, which are: remembering, understanding, applying, analysing, evaluating, and creating. This categorisation is meant to help educators to define learning goals, exercises, and questions accordingly. To stimulate reflection, the last three levels, analysing, evaluating, and creating, seem the most promising as they are also components of critical thinking (Paul, 2019), which aims at “judging in a reflective way what to do or what to believe” (Facione, 2000). Second, the Socratic method is a questioning method that is widely used in education to develop and promote critical thinking skills (Makhene, 2019; Ho et al., 2023; Dalim, 2022; Torabizadeh et al., 2018). The Socratic method consists of systematically asking questions that help to clarify concepts, improve understanding, and uncover gaps in knowledge (Paul, 2019). Further, it helps to discover own thoughts, to analyse and evaluate assumptions, information and inferences, and to arrive at judgements through own reasoning (Paul, 2019). To help teachers to explore and encourage their students’ thinking through guided questions, a taxonomy of Socratic questions has been developed, defining eight types of questions that probe different parts of thinking, such as assumptions, reasons, or implications (Paul, 2019).
@@
The Socratic questioning method is slowly gaining traction in the field of human-machine interaction. Based on the taxonomy of Socratic questions, a dataset of 110k questions that could be used in counselling and coaching is proposed, including a generation method of prompt-tuning a language model (Ang et al., 2023). The resulting questions fall into one of the six categories defined by Paul (2019): 1) clarification (“What do you mean by…?”), 2) probing assumptions (“Why do you assume…?”), 3) probing reasons (“How do you know that…?”), 4) probing implications (“If…, what is likely to happen?”), 5) probing alternatives (“What else should we consider?”), and 6) others that do not conform to the Socratic categories (Ang et al., 2023). In addition to this work, a concept of a virtual assistant that fosters moral enhancement through Socratic questions is proposed (Lara and Deckers, 2020), as well as chatbots based on language models that utilise the Socratic method for thought-provoking tutoring (Favero et al., 2024; Liu et al., 2024).
@@
While these works appear promising, there is currently no systematic approach to using questions to promote critical reflection in decision-making, especially in combination with decision-support systems. Given the link between critical reflection and reasoning and the associated positive effect on decision-making, we see potential for the use of questions to promote decision self-efficacy and cognitive engagement. We therefore aim to address this gap by transferring an existing taxonomy of Socratic questions (Paul, 2019, 2010) to the context of machine-assisted decision-making. In doing so, we also aim for a higher level of granularity as compared to Ang et al. (2023), who utilise parts of the Socratic questions taxonomy in unchanged form to generate counselling and coaching questions. With our proposed taxonomy we provide a systematic way to formulate relevant questions for critical reflection during decision-making, based on concepts of critical thinking and the Socratic questioning method.
@@ 3. Methodology
In the following, we will discuss how we constructed our question taxonomy, which we did in two steps: identifying possible questions, and identifying information for formulating questions. Overall, we synthesised three prior works, namely 1) a taxonomy of Socratic questions (Paul, 2019), 2) a question bank for the design of human-centred XAI (Liao et al., 2020, 2021; Liao and Varshney, 2022), and 3) a revision of Bloom’s taxonomy for educational objectives (Krathwohl, 2002; Forehand, 2010). The creation of the taxonomy resembles affinity diagramming with a deductive (top-down) approach, for which the Socratic questions functioned as a starting point and the XAI question bank helped us to translate the categories to the context of machine-assisted decision-making (Figure 1).
3.1. Phase 1: What to question?
In the first step, we identified relevant aspects of machine-assisted decision-making to which potential questions should relate, i.e.,what elements to reflect on. For this purpose, we used a taxonomy for Socratic questions based on critical thinking concepts (Paul, 2019, 2010). As Elder and Paul (2010) note:
“Whenever we think we think for a purpose within a point of view based on assumptions leading to implications and consequences. We use data, facts and experiences to make inferences and judgements based on concepts and theories to answer a questions or solve a problem.”
Accordingly, the authors categorise questions as follows into elements of thought or reasoning that allow to question the structure of thinking (Paul, 2019, 2010):
@@ 3.1. Phase 1: What to question?
While these elements of thought are helpful in general, the provided example questions (Table A1 in the Appendix) cannot be directly applied to the domain of decision-making (with DSSs). This is because the taxonomy of Socratic Questions is primarily intended for teachers and students in the classroom. We thus took the elements of thought as a starting point for our taxonomy and mapped them to the context of machine-assisted decision-making. The element Information, for example, can be mapped, as is also mentioned in the definition, to the dataset the DSS is trained on, as well as data of a specific (patient) case. To further specify relevant elements or aspects of machine-assisted decision-making, we used the XAI question bank, in which the authors cluster more than 50 common questions operators have towards a DSS into the following categories (Liao et al., 2020, 2021; Liao and Varshney, 2022):
@@
Questions from the XAI question bank are about the logic of the model, its performance, the reasons for a prediction, possible changes, the data, and the output.
@@
This leaves us with, on the one hand, the categorisation of questions based on the structure of thought, and on the other hand, the components of machine-assisted decision-making represented by clusters of common questions in the context of explainable AI (Figure 1). Through affinity diagramming we arrived at the following 10 elements: case information, relevance of data, dataset, causal structure of recommendation, alternatives to recommendation, assumptions of decision-maker, stakeholder preferences, consequences of decision, change intervention, and model behaviour.
To define our final question categories, we followed an iterative process. We presented 10 initial categories, including some sample questions, to one orthopaedic surgeon and one senior researcher in orthopaedics who confirmed the completeness of the derived question types. We then refined the categories, without adding or removing any, expanded details, added descriptions, and added sample questions to a separate table (Table A3).
3.2. Phase 2: How to enrich questions?
@@
After identifying what to question, the next step was to define how to derive and enrich specific questions of said categories. For this we again used the XAI question bank. For our purposes, however, we reversed the process. Instead of starting with questions from operators in order to arrive at suitable explanations, we use XAI techniques to extract information which then inform possible questions (column ‘Useful Information’ in Table 1). On the one hand, this means that (some) explanations can be (re-)formulated as questions (Danry et al., 2023), e.g., “Does outcome follow from information ?”. On the other hand, questions can relate to information extracted from explanations, e.g., “Is information the relevant factor to focus on?”.
@@
Finally, to capture the scope of questions, we used, albeit only marginally, a revised version of Bloom’s taxonomy (Krathwohl, 2002; Forehand, 2010; Anderson and Krathwohl, 2001). In Bloom’s taxonomy, questions are categorised hierarchically into six levels of cognitive processes, which are: remembering, understanding, applying, analysing, evaluating, and creating. The last three levels, analysing, evaluating, and creating, are also components of critical thinking (Paul, 2019). In order to effectively stimulate critical reflection, we thus hypothesise that questions should address these three higher-level cognitive abilities, where:
@@
Analysing means to break down information into parts and determining how these parts are connected, and identifying which parts are relevant or irrelevant. Verbs that relate to this cognitive process are, compare, criticize, differentiate, discriminate, deconstruct, inspect. For example, “What evidence is there for ?”.
@@
Evaluating means to judge information and its appropriateness, and to detect inconsistencies. Verbs that relate to this cognitive process are, assess, support, defend, validate, evaluate. For example, “How to justify ?”
@@
Creating means compiling information in a new way, or considering alternative hypotheses. Verbs that relate to this cognitive process are, create, develop, formulate. For example, “What could be done to minimise ?”
This distinction, with the associated verbs, is meant to help formulate concrete questions for specific use cases.
4. Taxonomy of Questions for Critical Reflection
@@
In the following, we will present our taxonomy, which consists of 10 question types or elements for reflection (Table 1). In order to demonstrate how to apply our taxonomy, we provide some sample questions from the medical and educational domain in Table A3 and Table A4 respectively.
@@
Questions that help evaluate whether the outcome follows from data to ensure that the causal structure of the model / recommendation is sound. | Feature Contribution (e.g., (Lundberg and Lee, 2017), (Ribeiro et al., 2016)), Counterfactuals |
@@
Questions that help to consider other possibilities ensuring that the larger solution space is considered. | Contextual Information (e.g., differential diagnosis or patient preferences (Q7)), Partial Dependence Plot, Counterfactuals, |
@@
Questions that elicit taken for granted assumptions to ensure that the decision-maker is aware of their reasons and can direct their actions accordingly. | General Questions, Confidence of decision-maker |
@@
Questions that help the decision-maker enquire and take into account the preferences or needs of the people concerned (e.g., patient), in order to ensure a better or broader understanding of the problem at hand. | General questions, Preference Elicitation |
@@
Questions that help elicit anticipation and forward-looking responsibility to ensure that (unintended) consequences are considered and mitigated. | Limitations of Model, FactSheets (Arnold et al., 2019), Model Cards (Mitchell et al., 2019) |
@@
Questions that support the exploration of feasible (smaller) interventions to make a desired outcome more likely. | Feature Perturbation, Counterfactuals (Karimi et al., 2021; Poyiadzi et al., 2020) |
@@
Q1. Questions can relate to the case information provided, which represents information in the form of data. Previous operations on the spine, for example, might be indicated by integers, but this value does not specify how long ago the surgery was performed and at which location of the spine it was carried out - both aspects that can influence the effectiveness of further operations. Questions can help the decision-maker to inspect and contextualise the data provided to give it semantic depth. In addition to ensuring data quality, questions can also relate to the reliability of the data provided to ensure that inconsistencies, such as measurement errors, are ruled out, and that no relevant data is missing. As such, Q1 helps the decision-maker get a more holistic picture of the case that goes beyond data-driven abstractions (van Baalen and Boon, 2015).
@@
Q2. Questions can relate to the relevance of data. XAI methods make it possible to extract the contribution of the various features to the outcome and rank them accordingly. It is possible to use that information to generate an evaluative question, like “How relevant is data ?”. By questioning the relevance of data points, the decision-maker is encouraged to contextualise the information, e.g., symptom, in a specific case. To do so, the decision-maker may wish to check and validate the machine recommendation by other means, e.g., by conducting additional tests. So while XAI can show which features have contributed to an outcome, domain knowledge is necessary to evaluate whether these features are important in a particular case. Questioning the relevance of the most contributing data point also raises the question whether there are other factors that are overlooked in the decision. In addition, considering that different XAI techniques can result in different rankings of the contributing effects of the individual features (Saarela and Jauhiainen, 2021), it is important for decision-makers not to take the extracted information provided by XAI at face value. The contradiction between two opposing feature ranks by two different XAI techniques could also lead to a question like “Is or more relevant?”.
@@
Q3. Questions can relate to the dataset on which the model was built. Cardiovascular diseases, for example, show different symptoms in men and women. Most of the collected data that is used to train DSS, however, stems from male patients, which is why DSS often lead to misdiagnosis in women (Li et al., 2023). The problem in this case is that certain populations or phenomena are overrepresented in data, which leads to a skewed data distribution (Ueda et al., 2024). Besides, data is often only a partial representation of the phenomena under investigation (Broussard, 2019; Binns et al., 2018), as it is impossible to quantify every aspect of the complexity of social life (Birhane, 2021). Questions can help to consider limitations of a dataset and take into account information that is not yet included and is (now) known to contribute to the outcome. So compared to Q2, which is more concerned with the question of what information is relevant for a single recommendation (local), Q3 asks whether the dataset contains relevant information (global).
@@
Q4. Questions can relate to the (apparent) causal structure between provided case information and resulted machine recommendation. Again, based on feature contribution it is possible to ask whether and how the machine recommendation follows from data point , for instance “Does diagnosis follow from symptom ?”. So although this question type merely reformulates a causal explanation (Danry et al., 2023), the aim is to help the decision-maker to analyse and evaluate information.
@@
Q5. Questions can help identify and consider alternatives to the recommendation. This calls into question the solution of the DSS and thus, similar to Q4, the appropriateness of a recommendation. Different disease have common symptoms, so the physician wants to make sure that another possible diagnosis is not overlooked. For this, a differential diagnosis can be helpful (Mamede and Schmidt, 2023). At the same time, it can be possible to treat the same condition in different ways. Chronic lower back pain, for example, might be treated by surgical or non-surgical interventions (van Hooff et al., 2014), where the feasibility of each option can depend on constraints like available resources or the patient’s preferences (Q7). For example, if the purpose of a DSS is to optimise for full recovery, it might recommend surgery that involves a long stay in hospital; another purpose might be to promote the well-being of the patient, who may prefer to spend as little time as possible in hospital, so that the more suitable treatment could be a non-surgical intervention. Q5 supports the decision-maker in developing alternative hypotheses and courses of action.
@@
Q6. Questions can relate to the assumptions of the decision-maker. On the one hand, these questions can help to elicit tacit knowledge and taken for granted assumptions, which might need to be re-evaluated on a case to case basis. This could prevent the decision-maker from jumping to conclusions. On the other hand, questions can help the decision-maker to increase their confidence and self-efficacy as they allow them to arrive at their own decision. Moreover, assumptions of the decision-maker could also be compared to the machine recommendation.
@@
Q7. Questions can relate to stakeholder preferences. Especially in decision contexts in which decision-maker and decision-subject are different persons, such as physician and patient, it is important to take into account different perspectives. Ideally, enquiring about the preferences of stakeholders (patient) is part of the decision-maker’s (physician) routine practice, yet these questions can function as a form of safeguard. In addition, in case the decision-maker is the decision-subject, questions can also help elicit preferences of the decision-maker themselves.
@@
Q8. Questions can help to anticipate (unintended) consequences of a decision. While the decision-maker has ideally weighed up the implications, these questions can promote forward-looking responsibility of the decision-maker. Furthermore, with insights from the dataset (Q3) attention could be drawn to certain limitations of a DSS. Compared to Q3, the focus of Q8 is on assessing the impact of a recommendation, and supporting the decision-maker in considering appropriate mitigation strategies.
@@
Q9. Questions can relate to change interventions in order to further investigate the space of different alternatives. By creating hypothetical scenarios, the feasibility of desired outcomes can be investigated. In doing so, the decision-maker is encouraged to consider prior and smaller interventions that could increase the expected effectiveness of a desired treatment. To create such question, feature perturbation and counterfactual explanations can be helpful (Karimi et al., 2021; Poyiadzi et al., 2020). In contrast to counterfactual explanations, which provide information about what the outcome would have been if the data had been like this, these questions are about whether it is possible to change certain data points in order to make a particular outcome more likely. For example, a counterfactual explanation would be “If the salary were higher, the loan would have been accepted”, whereas the question turns into “Is it possible increase the salary, so that the probability of an accepted loan increases?”. Although these questions ideally lead to more actionable insights, they might imply unfeasible or unattainable changes and actions, e.g., changing the age of a patient. These hypothetical scenarios can nevertheless provide insights into inbuilt associations and decision boundaries of the DSS.
@@
Q10. Questions can relate to the model behaviour. For example, if a DSS makes a decision based on age, with a threshold of 50 years, the decision-maker may want to re-evaluate the decision, if the patient is close to that threshold, e.g., 48 years old. These questions might also relate to the relevance of data points (Q1), as single data points often function as proxies for other characteristics. Age, for example, might be an indication for how well the patient will recover from surgery, yet physical fitness might also play a role. This means, focusing only on single data points could be misleading. As such, these questions can also give the decision-maker a better understanding of the causal structure built into the model (Q4).
5. Evaluation of Our Taxonomy

### v3

@@ 2.5. Types of Questions
A common technique to facilitate reflection is to ask questions (Osmond and Darlington, 2005). Questions encourage independent thinking, which can increase confidence in one’s own decision-making, i.e., decision self-efficacy. This in turn can affect the motivation and well-being of the decision-maker (Ryan and Deci, 2000). Findings from educational research suggest that motivation functions as an antecedent to cognitive engagement and also increases it (Blumenfeld et al., 2006; Rotgans and Schmidt, 2011; Singh et al., 2022). Questions thus seem to promote reflection, which can increase decision self-efficacy and motivation, which in turn can increase cognitive engagement.
Given the breadth of possible questions, we want to point out two works from the field of education that categorise questions in different ways and which we also use for our purposes, namely a taxonomy of Socratic questions (Paul, 2019), and Bloom’s taxonomy for educational objectives (Anderson and Krathwohl, 2001; Krathwohl, 2002).
@@
First, Socratic questioning is widely used in education to develop and promote critical thinking skills (Torabizadeh et al., 2018; Makhene, 2019; Dalim, 2022; Ho et al., 2023). It consists of systematically asking questions that help to clarify concepts, improve understanding, and uncover gaps in knowledge (Paul, 2019). Accordingly, Socratic questioning helps to discover one’s thoughts, to analyse and evaluate assumptions, information and inferences, and to arrive at judgements through own reasoning. To support teachers in the classroom, a taxonomy of Socratic questions has been developed, defining eight types of questions that probe different parts of thinking, such as assumptions, reasons, or implications (Paul, 2019).
Second, Bloom’s taxonomy categorises educational learning objectives hierarchically into six levels of cognitive processes, which are: remembering, understanding, applying, analysing, evaluating, and creating. The last three levels, analysing, evaluating, and creating, are also components of critical thinking (Paul, 2019). This categorisation will be helpful when formulating questions that address different cognitive processes.
@@
In the context of human-machine interaction, the Socratic questioning method is slowly gaining traction. Based on the taxonomy of Socratic questions, a dataset of 110k questions that could be used in counselling and coaching is proposed, including a generation method of prompt-tuning a language model (Ang et al., 2023). The resulting questions fall into one of the six categories defined by Paul (2019): 1) clarification (“What do you mean by…?”), 2) probing assumptions (“Why do you assume…?”), 3) probing reasons (“How do you know that…?”), 4) probing implications (“If…, what is likely to happen?”), 5) probing alternatives (“What else should we consider?”), and 6) others that do not conform to the Socratic categories (Ang et al., 2023). In addition, a virtual assistant that fosters moral enhancement through Socratic questions is proposed (Lara and Deckers, 2020), as well as chatbots based on language models that utilise the Socratic method for thought-provoking tutoring (Favero et al., 2024; Liu et al., 2024).
@@
However, there is currently no systematic approach to the use of (Socratic) questions to promote critical reflection in decision-making, particularly in conjunction with decision-support systems. Given the link between critical reflection and reasoning and the associated positive effect on decision-making, we see potential for the use of questions to promote decision self-efficacy and cognitive engagement. We therefore aim to address this gap by adapting the taxonomy of Socratic questions (Paul, 2010, 2019) to the context of machine-assisted decision-making. In doing so, we also aim for a higher level of granularity as compared to Ang et al. (2023), who utilise parts of the Socratic questions taxonomy in unchanged form to generate counselling and coaching questions.
@@ 3. Methodology
Our taxonomy is based on the taxonomy of Socratic questions (Paul, 2019), the XAI question bank (Liao and Varshney, 2022), and a revision of Bloom’s taxonomy (Anderson and Krathwohl, 2001) discussed in the previous section. To construct our taxonomy we proceeded in two steps: 1) identifying elements that questions can relate to, and 2) identifying information for formulating and enriching questions. In order to identify elements (1), we synthesised the taxonomy of Socratic questions with the XAI question bank through a procedure similar to affinity diagramming. In order to help formulate and enrich questions (2), we reversed the mapping of the XAI question bank to repurpose information derived from explanations. In addition, we used the cognitive processes identified in Bloom’s taxonomy to determine the scope of questions.
3.1. Procedure
3.1.1. Phase 1: Identifying Elements for Reflection
First, we identified relevant aspects of machine-assisted decision-making to which potential questions should relate, i.e., what to reflect on. In a deductive (top-down) approach, the categories from the taxonomy of Socratic questions served as a starting point. As Elder and Paul (2010) note:
“Whenever we think we think for a purpose within a point of view based on assumptions leading to implications and consequences. We use data, facts and experiences to make inferences and judgements based on concepts and theories to answer a question or solve a problem.”
Based on this structure of thinking, the authors categorise questions into elements of thought or reasoning as follows:
@@ 3.1.1. Phase 1: Identifying Elements for Reflection
While these elements of thought, including the provided example questions (Table A), are helpful in general, they cannot be directly applied to the domain of decision-making (with DSSs). This is because the taxonomy of Socratic Questions is primarily intended for teachers and students in the classroom.
We thus took the elements of thought as a starting point for our taxonomy and mapped them to the context of machine-assisted decision-making. For this mapping and to identify relevant aspects of machine-assisted decision-making, we used the XAI question bank. The question bank contains more than 50 common questions that operators have towards a DSS, categorised as follows (Liao and Varshney, 2022):
@@
Overall, questions from the XAI question bank relate to the logic of the model, its performance, the reasons for a prediction, possible changes, the data, and the output.
@@
This leaves us with the categorisation of questions based on thought structure on the one hand, and the components of machine-assisted decision-making represented by clusters of common questions in the context of explainable AI on the other (Figure 1). The element Information, for example, can be mapped, as is also mentioned in the definition, to Data the DSS is trained on, as well as (input) data of a specific case. Through affinity diagramming we arrived at the following 10 elements, which are shown in the middle column of Figure 1: case information, relevance of data, dataset, causal structure of recommendation, alternatives to recommendation, assumptions of decision-maker, stakeholder preferences, consequences of decision, change intervention, and model behaviour.
A sankey plot showing the connections between the elements of our taxonomy and the taxonomy for Socratic questions and the XAI question bank.
3.1.2. Phase 2: Formulating Questions
@@
After identifying what to question, the next step was to specify how to formulate and enrich specific questions. For this, we again used the XAI question bank, in particular the mapping of common questions to XAI techniques. Instead of starting with questions from operators in order to arrive at suitable explanations, however, we reversed the process. We moved from explanations to questions, in order to extract information using XAI techniques, which then flows into possible questions. On the one hand, this means that (some) explanations can be (re-)formulated as questions, e.g., “Does outcome follow from information ?” (Danry et al., 2023). On the other hand, questions can relate to information extracted from explanations, e.g., “Is information a relevant factor?”.
@@
Finally, to capture the scope of questions, we used, albeit only marginally, a revised version of Bloom’s taxonomy. In Bloom’s taxonomy, questions are categorised into levels of cognitive processes to support the formulation of learning objectives. Three of these categories or cognitive abilities are components of critical thinking (Paul, 2019), namely, analysing, evaluating, and creating, where (Anderson and Krathwohl, 2001):
@@
Analysing means to break down information into parts and determining how these parts are connected, and identifying which parts are relevant or irrelevant. Verbs that relate to this cognitive process are: compare, criticize, differentiate, discriminate, deconstruct, inspect. For example, “What evidence is there for ?”.
@@
Evaluating means to judge information and its appropriateness, and to detect inconsistencies. Verbs that relate to this cognitive process are: assess, support, defend, validate, evaluate. For example, “How to justify ?”
@@
Creating means compiling information in a new way, or considering alternative hypotheses. Verbs that relate to this cognitive process are: create, develop, formulate. For example, “What could be done to minimise ?”
Based on this cognitive distinction, we suggest that questions should address these higher-level cognitive processes in order to effectively stimulate critical reflection.
3.2. Expert Feedback
While one author created the initial taxonomy, we discussed the categories with their overlaps and boundaries among the authors throughout the process. Besides, we created the taxonomy in an iterative way. Once we had a first version, we sent 10 initial categories including some sample questions to one orthopaedic surgeon and one senior researcher in orthopaedics, who both work at different hospitals and are part of an ongoing collaboration. To get some early feedback, we asked them if they saw any missing question types. They confirmed the completeness of the derived question types and mentioned that it seems we “covered all the aspects” and that they “don’t see any obvious questions missing”. We then refined the categories, without adding or removing any, expanded details, added descriptions, and added sample questions to a separate table (Table C). We presented the final taxonomy again to the same surgeon and researcher (section 5).
4. Taxonomy of Questions
@@
Our taxonomy consists of 10 question types or elements for reflection (Table 1). In order to demonstrate how to apply our taxonomy, we provide some sample questions from the medical (Table C) and educational domain (Table D). A description of the individual question type follows, focussing on the medical field, where a physician uses a DSS to make a decision about a patient’s treatment.
@@
Questions that help evaluate whether the outcome follows from data to ensure that the causal structure of the model / recommendation is sound. | Feature Contribution (Lundberg and Lee, 2017; Ribeiro et al., 2016), Counterfactuals (Karimi et al., 2021; Poyiadzi et al., 2020) |
@@
Questions that help to consider other possibilities ensuring that the larger solution space is considered. | Contextual information (e.g., differential diagnosis or patient preferences (Q7)), Partial Dependence Plot, Counterfactuals |
@@
Questions that elicit taken-for-granted assumptions to ensure that the decision-maker is aware of their reasons and possible cognitive biases. | General questions, Cognitive biases, Confidence of decision-maker |
@@
Questions that help the decision-maker enquire and take into account the preferences or needs of the people concerned (e.g., patient), in order to ensure a better or broader understanding of the problem at hand. | General questions, Preference elicitation, Values |
@@
Questions that help elicit anticipation and forward-looking responsibility to ensure that (unintended) consequences and trade-offs are considered and mitigated. | Limitations of Model, FactSheets (Arnold et al., 2019), Model Cards (Mitchell et al., 2019) |
@@
Questions that support the exploration of feasible (smaller) interventions to make a desired outcome more likely. | Feature Perturbation, Counterfactuals |
@@
Q1. Questions can relate to the case information provided, which represents information in the form of data. Previous operations on the spine, for example, might be indicated by integers. This value, however, does not indicate how long ago the operation was performed and at which location of the spine it was carried out, both of which can influence the effectiveness of further operations. Questions can help the decision-maker to inspect and contextualise the data provided to give it semantic depth. To ensure data quality, questions can also relate to the reliability of the data provided to rule out inconsistencies, such as measurement errors, and to ensure that no relevant data is missing. As such, Q1 helps the decision-maker get a more holistic picture of the case (van Baalen and Boon, 2015).
@@
Q2. Questions can relate to the relevance of data. In doing so, questions can encourage the decision-maker to contextualise the information, e.g., symptom, in a specific case, which can lead to validating the machine recommendation by other means, e.g., by conducting additional tests. So while XAI techniques, such as feature importance, can show which features have contributed to an outcome, domain knowledge is necessary to evaluate whether these data points are important in a particular case. It is possible to use the information extracted with XAI to generate an evaluative question, like “How relevant is data ?”. The question of the relevance of the most important data point(s) also raises the question of whether there are other factors that could be overlooked in the decision. As different XAI techniques can result in different rankings of the contributing effects of the individual features (Saarela and Jauhiainen, 2021), it is important for decision-makers not to take the information obtained through XAI at face value. The discrepancies between two opposing feature ranks from two different XAI methods could also lead to a question like “Is or more relevant?”.
@@
Q3. Questions can relate to the dataset on which the model was built. Cardiovascular disease, for example, has different symptoms in men and women, yet most of the data for training DSSs stem from male patients, which is why DSSs often misdiagnose women (Li et al., 2023). Certain populations or phenomena are often overrepresented in data, which leads to a skewed data distribution (Ueda et al., 2024). Accordingly, a reflective question could be raised if there are large differences between the case information and the dataset. Besides, data is often only a partial representation of the phenomena under investigation (Broussard, 2019; Binns et al., 2018), as it is impossible to quantify every aspect of the complexity of social life (Birhane, 2021). Questions can help to consider limitations of a dataset and take into account information that is not yet included and is (now) known to contribute to the outcome. So compared to Q2, which is more concerned with the question of what information is relevant for a single recommendation (local), Q3 asks whether the dataset contains relevant information (global).
@@
Q4. Questions can relate to the (apparent) causal structure between provided case information and resulted machine recommendation. Again, based on feature contribution it is possible to ask whether and how the machine recommendation follows from data point , for instance “Does diagnosis follow from symptom ?”. So although this question type merely reformulates a causal explanation (Danry et al., 2023), the aim is to help the decision-maker to scrutinise the information.
@@
Q5. Questions can help identify and consider alternatives to the recommendation. This calls into question the appropriateness of a recommendation. It might be possible to treat the same condition in different ways, where the feasibility of each option depends on various constraints like available resources or the patient’s preferences (Q7). For example, one purpose might be full recovery, for which a DSS might recommend surgery that involves a long hospital stay; another purpose might be to promote the patient’s well-being, who might prefer to spend as little time as possible in hospital, so that the more appropriate treatment might be a non-surgical intervention. At the same time, as different disease have common symptoms, the physician wants to ensure that another possible diagnosis is not overlooked, e.g., through a differential diagnosis (Mamede and Schmidt, 2023).
@@
Q6. Questions can relate to the assumptions of the decision-maker. On the one hand, these questions can prevent the decision-maker from jumping to conclusions by helping to uncover tacit knowledge, taken for granted assumptions, and cognitive biases and re-evaluate them case by case. On the other hand, questions can help to increase the confidence and self-efficacy of decision-makers as they allow them to make their own decision.
@@
Q7. Questions can relate to stakeholder preferences. If the decision-maker is not the decision-subject, as is the case with physicians and patients, it is important to consider the different perspectives. Ideally, enquiring about the preferences of stakeholders is part of the decision-maker’s routine practice, yet these questions can function as a form of safeguard. If the decision-maker is the decision-subject, questions can help elicit preferences and values of the decision-maker themselves.
@@
Q8. Questions can help anticipate (unintended) consequences of a decision. While the decision-maker has ideally weighed up the implications, these questions can promote forward-looking responsibility of the decision-maker. Furthermore, with insights from the dataset (Q3) attention could be drawn to certain limitations of a DSS. Compared to Q3, Q8 focuses on assessing the impact of a recommendation, weighing trade-offs, and supporting the decision-maker in considering appropriate mitigation strategies.
@@
Q9. Questions can relate to change interventions in order to further investigate the feasibility of desired alternatives. For such questions, hypothetical scenarios can be created by feature perturbation and counterfactual explanations. Unlike counterfactual explanations, which provide information about what the outcome would have been if the data had been like this, these questions are about whether it is possible to change certain data points to make a particular outcome more likely. For example, a question might help the physician to consider prior and smaller interventions such as smoking cessation, which could increase the expected effectiveness of a desired treatment such as surgery. Although these questions ideally lead to more actionable insights, they might imply unfeasible or unattainable changes and actions, e.g., changing the age of a patient. These hypothetical scenarios can nevertheless provide insights into inbuilt associations and decision boundaries of the DSS.
@@
Q10. Questions can relate to the model behaviour. These questions help to clarify the assumptions built into the DSS, which may be based on the data set (Q3) or other assumptions made during model development. As such, these questions can also provide the decision-maker with a better understanding of the workings and (apparent) causal structure built into the model (Q4). For example, if a DSS makes a decision based on age with a threshold of 50 years, the decision-maker may want to re-evaluate the decision, if the patient is close to that threshold, such as 48 years old.
5. Evaluation
To formatively evaluate our question taxonomy, we presented a first version with sample questions and the final version in two phases to the same orthopaedic surgeon and orthopaedic researcher. In individual online meetings of 30 minutes each, which were not recorded or transcribed, we asked them whether the taxonomy and the derived questions could help practitioners, whether the considerations associated with these questions already correspond to common practice, whether certain questions might be more relevant than others, and whether they could provide example questions or example cases.